OpenStudy (anonymous):

one simple integral! x/sq.rt(1-x^4)??

OpenStudy (anonymous):

\[\int\limits_{\frac{ -1 }{ \sqrt{2} }}^{\frac{ 1 }{ \sqrt{2} }}\frac{ x }{ \sqrt{1-x^4} }\]

OpenStudy (experimentx):

let x^2 = y

OpenStudy (experimentx):

then use trig subs.

OpenStudy (anonymous):

i did. u=x^2 du=2xdx \[\frac{ 1 }{ 2 }\int\limits_{-1/2}^{1/2}\frac{ du }{ \sqrt{1-u^2} }=\sin^{-1}u \]

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

Method 1: Note that \(\large\displaystyle \int_{-1/\sqrt{2}}^{1/\sqrt{2}} \frac{x}{\sqrt{1-x^4}}\,dx = \int_{-1/\sqrt{2}}^{1/\sqrt{2}} \frac{x}{\sqrt{1-(x^2)^2}}\,dx\). Let \(\large u=x^2\) and recall that \(\large\displaystyle\int \frac{1}{\sqrt{1-y^2}}\,dy = \arcsin y + C\) Method 2: Note that \(\large \dfrac{x}{\sqrt{1-x^4}}\) is an odd function. Do you recall what the integral of an odd function over a symmetric interval \(\large [-a,a]\) is?

OpenStudy (anonymous):

@emilyoo the lower limit of your new integral isn't correct. Note that \(\large \left(-\dfrac{1}{\sqrt{2}}\right)^2 = \dfrac{1}{2}\).

OpenStudy (anonymous):

oh i see i made a mistake

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

No problem. So what did you get in the end? :-)

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Good. :-)